The circuit below is "Differential amplifier" The waveform of the input signal is shown below. The input signal is V_{in(rms)} =250mV , f_{in}=20kHz

(1) An expression of the gain of a closed loop on a circuit R_{1},R_{2},R_{_{f}},R_{g}
Use to induce. (formerly op-amp)
The analysis method can be used. KVL and KCL can be utilized properly.)

(2)When R_{1}=R_{2},R_{_{f}}=R_{g} set the resistance value so that the closed loop gain is 2.5 and the output synchronized with the input signal.
Sketch the force signal (exactly mark the maximum and minimum values and cycles of the output signal on the output signal waveform).
Answer #1

Ri ए Uza Rg The above circuit is a difference Amplifier which means output voltage is directly propastional to difference ofNow the circuit is £ Ri > 쿄 Rg a ارا Applying KCL at ne rode FIB + If For an ideal op-Amp 18 = 15 co I1 = If u-va = va-vo R Rdue to y V- Ri o-vo, RE Voi = -RE RI Y Now we have to calculate output voltage due to non Investing terminal at the same timeof Ri Rz Vo=11HRE RI then vo=114 REV + (14RE Ritky). 112 etky) - Rey 4 Rp = Rz, Rf = Rg -len Vo = lethel Petre) k een -V gai